# Bridged Semantic Alignment for Zero-shot 3D Medical Image Diagnosis11footnotemark: 1

### v1

@@ Abstract
3D medical images such as Computed tomography (CT) are widely used in clinical practice, offering a great potential for automatic diagnosis. Supervised learning-based approaches have achieved significant progress but rely heavily on extensive manual annotations, limited by the availability of training data and the diversity of abnormality types. Vision-language alignment (VLA) offers a promising alternative by enabling zero-shot learning without additional annotations. However, we empirically discover that the visual and textural embeddings after alignment endeavors from existing VLA methods form two well-separated clusters, presenting a wide gap to be bridged. To bridge this gap, we propose a Bridged Semantic Alignment (BrgSA) framework. First, we utilize a large language model to perform semantic summarization of reports, extracting high-level semantic information. Second, we design a Cross-Modal Knowledge Interaction (CMKI) module that leverages a cross-modal knowledge bank as a semantic bridge, facilitating interaction between the two modalities, narrowing the gap, and improving their alignment. To comprehensively evaluate our method, we construct a benchmark dataset that includes 15 underrepresented abnormalities as well as utilize two existing benchmark datasets. Experimental results demonstrate that BrgSA achieves state-of-the-art performances on both public benchmark datasets and our custom-labeled dataset, with significant improvements in zero-shot diagnosis of underrepresented abnormalities.
keywords:
@@
journal: Medical Image Analysis
Fig. 1: UMAP visualization of features. Cosine similarity is used to evaluate the alignment between image and text features. The text features are generated using generic descriptive texts to ensure that all images can be matched to all texts. (a) Features generated using pretrained weights without vision-language alignment, where image (in blue) and text features (in red) remain unaligned. (b) Features after vision-language alignment using 3D CLIP, showing improved alignment but with noticeable modality gaps. (c) Features after vision-language alignment using BrgSA framework, where the CMKB features (in green) serve as a bridge to reduce the modality gap and further enhance feature alignment.
1 Introduction
@@
Computed Tomography (CT) is a widely utilized non-invasive diagnostic tool in clinical practice, playing a crucial role in detecting various abnormalities [13, 31, 3]. With the rapid development of artificial intelligence, significant progress has been made in leveraging CT images for automated abnormality diagnosis [39]. Supervised learning methods have achieved impressive performance in tasks such as disease classification and detection [16, 45, 52, 25, 51, 33, 43, 30]. However, supervised learning methods heavily rely on large-scale, manually annotated datasets, which is both time-consuming and labor-intensive [17, 12]. Moreover, the reliance on extensive annotations limits the diversity of available training data, particularly for rare diseases, where labeled samples are are exceedingly scarce [40]. This constraint hampers the scalability of supervised learning methods and diminishes their ability to generalize across a wide spectrum of diseases [49].
@@
Recently, zero-shot learning (ZSL) based on vision-language alignment has emerged as a promising alternative to traditional supervised learning methods [28, 55]. By eliminating the need for extensive manual annotations, ZSL offers a more efficient path to diagnose a diverse range of abnormalities. For instance, CLIP [34] leverages contrastive learning on large-scale image-text datasets to enable effective vision-language alignment and perform zero-shot classification without extensive manual annotations [56]. In the medical domain, several studies have demonstrated the potential of CLIP for zero-shot abnormality diagnosis on 3D CT volumes [15, 6, 4, 2]. However, our experiments reveal that directly applying CLIP to 3D medical images remains inadequate in vision-language alignment, as the embedding space exhibits a significant modality gap. Specifically, Fig. 1(a) shows the feature space when only pretrained weights are used without introducing vision-language alignment; a clear gap appears between the two clusters of image and text features. The feature alignment score (FAS), which is a metric that quantifies the degree of vision-language alignment and is mathematically defined later in Section 5, is only 0.08. Note that the FAS ranges from 0 to 1, where 0 indicates no alignment between features and 1 signifies complete alignment. Fig. 1(b) represents the feature space after applying CLIP for 3D vision-language alignment. It is evident that, after alignment, the gap between image and text features appears more narrow than before alignment (FAS is improved to 0.37 from 0.08) but remains pronounced, fundamentally limiting the model’s ability to fully exploit zero-shot abnormality diagnosis on 3D CT volumes. These findings are consistent with previous studies in the natural domain that have identified a pronounced modality gap within the CLIP embedding space, hindering alignment performance [26, 42, 36, 11]. This highlights the necessity of addressing the modality gap in CLIP embeddings to achieve better alignment in 3D medical images.
@@
In this paper, we propose a Bridged Semantic Alignment (BrgSA) framework, which aims to bridge the modality gap and facilitate effective vision-language alignment. Our approach consists of two modules: semantic summarization and cross-modal knowledge interaction (CMKI). First, we utilize a large language model (LLM), which has powerful semantic understanding capabilities, to summarize clinical reports and extract critical information. This significantly reduces the learning difficulty on the textual side and produces high-level semantic features. Next, we propose a CMKI module to supplement the conventional CLIP alignment, where a cross-modal knowledge bank (CMKB) acts as a semantic bridge between image and text features. By projecting image and text features into the shared latent space of the CMKB, it reduces the discrepancy between modalities and simultaneously preserves their unique characteristics. This enables an implicit alignment approach that does not rely on paired image-text data. Finally, we employ contrastive learning to compare positive and negative pairs, achieving explicit vision-language alignment. Fig. 1(c) visualizes CMKB serving as a bridge between image and text features, effectively narrowing the modality gap. Also, the FAS is further improved to 0.50, indicating a better alignment.
@@
To explore the potential of our vision-language alignment approach for zero-shot diagnosis, we expand the original 18 abnormality labels on CT-RATE [15] with 15 additional labels. The new benchmark, referred to as “CT-RATE-LT,” focuses on underrepresented abnormalities with limited occurrences in the dataset, providing a challenging yet valuable testing ground for evaluating diagnostic methods. Experimental results on CT-RATE-LT demonstrate that our method achieves excellent performance in diagnosing underrepresented abnormalities, showing significant improvements compared to the existing state-of-the-art (SOTA) method (AUC: 76.9 vs. 85.6). Furthermore, we evaluate our method on the benchmark CT-RATE and RAD-ChestCT [10] datasets, achieving SOTA performances in zero-shot abnormality diagnosis, with an AUC of 79.2 and 70.0 on internal and external validation, respectively. In addition to its diagnostic capabilities, our method significantly outperforms the SOTA in the report-to-volume retrieval task (Recall@10: 5.0 vs. 10.1), demonstrating effective semantic alignment between image and text features.
@@
Fig. 2: Illustration of the proposed BrgSA network, which integrates semantic summarization and cross-modal knowledge interaction (CMKI). First, we leverage a large language model (LLM) to summarize the report, generating outputs in a fixed template. These summarized reports, along with the original reports, serve as the textual inputs. Then, image and text features are extracted by respective encoders and fed into CMKI module to obtain interaction features. Finally, the interaction features are constrained using an MSE loss, while alignment optimization is achieved via an InfoNCE loss.
@@ 2.1 Medical Vision-language Pretraining
Existing medical vision-language pretraining (VLP) research is predominantly focused on 2D imaging. Open-source datasets like MIMIC-CXR [20] have provided a solid foundation for the advancement of 2D VLP, while PadChest [5], which encompasses 193 abnormality categories, has established an effective benchmark for rare disease diagnosis. GLoRIA [18] enhances the ability to capture associations between images and text by integrating global and local feature alignment. CheXZero [41] leverages a CLIP model pre-trained on natural data, achieving stable improvements in medical VLP performance. MedCLIP [46] employs unpaired medical images and text with a semantic matching loss to mitigate false negatives. Xplainer [32] introduces an explainable zero-shot diagnosis framework based on observation-driven contrastive learning. MedKLIP [47] and KAD [54] incorporate domain-specific medical knowledge to improve VLP performance in CXR diagnosis tasks. CARZero [22] advances cross-modal alignment by leveraging cross-attention mechanisms to address the complex relationships between visual and textual modalities. These advancements have established 2D VLP as a robust framework for various abnormalities diagnosis.
@@ 2.2 Multi-Modality Interaction
Cross-modal interaction has been extensively explored to bridge the gaps between different modalities, such as vision and language modalities. Transformer-based approaches like ViLBERT [29] employs dual-stream transformers with co-attentional layers for modality-specific processing, while ViLT [21] uses a unified transformer to jointly encode visual patches and textual embeddings, enabling more efficient cross-modal interaction. MGCA [44] introduces cross-attention mechanisms to achieve cross-modal interaction, aligning medical image and text representations at multiple levels. MPMA [53] enhances cross-modal interaction by integrating image-text reconstruction with a global and local alignment mechanism, enabling richer semantic representation learning. However, these approaches primarily emphasize explicit alignment mechanisms, often overlooking the potential of leveraging shared semantic spaces to guide cross-modal learning effectively.
Dictionary learning has emerged as a powerful technique for enhancing cross-modal alignment by constructing shared representational spaces. Deng et al. [8] aligns features via a shared label space to improve cross-modal retrieval. HCDDL [23] builds hierarchical semantic embeddings by dictionary learning for fine-grained alignment. UNIMO-2 [24] leverages grounded semantic spaces to bridge visual and textual modalities, effectively addressing misalignment issues. Methods such as DetCLIP [48] and SOHO [19] further exploit trainable dictionaries to enrich visual-semantic representations, achieving SOTA results in zero-shot and vision-language tasks. CP-CLIP [50] employs the Core-Periphery principle to facilitate cross-modal interaction by structuring shared semantics and modality-specific features within a unified latent space. Despite these advancements, existing methods often fail to fully exploit the potential of shared representational spaces in improving vision-language alignment. To address this limitation, we propose a cross-modal knowledge interaction module that leverages shared representational spaces to achieve semantic-guided image learning without risking cross-modal information leakage.
@@ 3 Method
In this section, we describe our proposed BrgSA framework for zero-shot classification. As illustrated in Fig. 2, BrgSA consist of semantic summarization and CMKI module. More details are introduced in the following.
@@ 3.1 Semantic Summarization
To simplify learning on the textual side, we introduce semantic summarization. In the context of medical reports, textual data often contains lengthy and complex descriptions, including both critical abnormality-related information and a significant amount of irrelevant details. Such data characteristics increase the difficulty of extracting and aligning core semantic information with image features. We leverage the powerful semantic understanding and summarization capabilities of LLM to produce semantic summaries of medical reports, which are much easier to understand. As illustrated in Fig. 3, a prompt is designed to guide the LLM in extracting key abnormality-related information and generating outputs in a fixed template: “There is [abnormality].” Compared to directly processing the original text, this template-based summarization significantly reduces the difficulty of learning on the textual side. Furthermore, by combining the original reports with the LLM-generated summaries as input, the semantic richness of the textual data is preserved.
Fig. 3: Prompt for LLM used in semantic summarization for reports.
@@ 3.2 Feature Extraction
Assume that the training dataset contains $N$ samples denoted as $S_{\text{train}}=\left\{\left(I_{1},R_{1}\right),\ldots,\left(I_{N},R_{N}\right)\right\}$, where $I_{i}\in\mathbb{R}^{H\times W\times D}$ represents a 3D CT volume and $R_{i}$ represents its corresponding medical report. Here, $H,W,D$ denote the height, width, and depth of the 3D CT volumes, respectively. As illustrated in Fig. 2, we introduce the components of our architecture for feature extraction, including an image encoder $\mathbf{\Phi}_{\text{I}}$ for CT images and a text encoder $\mathbf{\Phi}_{\text{R}}$ for medical reports.
Image Encoder The image encoder is designed to extract global features from 3D CT images, as shown in Eq. 1.
@@
where $V_{i}\in\mathbb{R}^{d}$ represents the global features of the image, and $d$ denotes the feature dimension. In our experiments, we adopt a ViT-base [9] model as the image encoder. The [CLS] token is used to represent the global feature for 3D image.
Text Encoder The text encoder is used to extract global features from the medical report, as shown in Eq. 2.
@@
where $T_{i}\in\mathbb{R}^{d}$ represents the global features of the medical report. We use PubMedBERT [14] as the text encoder, which is well-suited for medical-related tasks due to its pretraining on biomedical literature. We use global average pooling for all non-[PAD] tokens as the global features of the text. In our experiment, the output feature dimensions of both the image and text encoders are unified to $d$.
3.2.1 Cross-Modal Knowledge Interaction
@@
In the CMKI module, we introduce a CMKB as a bridge for image-text interaction. The CMKB is initialized as a set of learnable embeddings:
@@
where $K$ represents the number of basis vectors.
@@
Using the attention weights $\mathbf{Z}_{i}^{V}$ and $\mathbf{Z}_{i}^{T}$, the original image and text features are reconstructed by taking a weighted sum of the CMKB basis vectors:
@@
where $\alpha$, $\beta$, and $\gamma$ are hyperparameters controlling the contributions of the reconstruction loss ($\ell_{\text{MSE}}$), the InfoNCE loss ($\ell_{\text{INFO}}$), and the InfoNCE loss on reconstructed features ($\ell_{\text{INFO-R}}$), respectively.
Fig. 4: Histogram of abnormality frequencies for CT-RATE-LT.
Fig. 5: Mapping of 27 abnormalities from RAD-ChestCT to 18 abnormalities in CT-RATE. The abnormalities from CT-RATE are denoted in blue font, whereas the abnormalities from RAD-ChestCT are denoted in black font.

### v2

@@ Abstract
3D medical images such as computed tomography are widely used in clinical practice, offering a great potential for automatic diagnosis. Supervised learning-based approaches have achieved significant progress but rely heavily on extensive manual annotations, limited by the availability of training data and the diversity of abnormality types. Vision-language alignment (VLA) offers a promising alternative by enabling zero-shot learning without additional annotations. However, we empirically discover that the visual and textural embeddings after alignment endeavors from existing VLA methods form two well-separated clusters, presenting a wide gap to be bridged. To bridge this gap, we propose a Bridged Semantic Alignment (BrgSA) framework. First, we utilize a large language model to perform semantic summarization of reports, extracting high-level semantic information. Second, we design a Cross-Modal Knowledge Interaction module that leverages a cross-modal knowledge bank as a semantic bridge, facilitating interaction between the two modalities, narrowing the gap, and improving their alignment. To comprehensively evaluate our method, we construct a benchmark dataset that includes 15 underrepresented abnormalities as well as utilize two existing benchmark datasets. Experimental results demonstrate that BrgSA achieves state-of-the-art performances on both public benchmark datasets and our custom-labeled dataset, with significant improvements in zero-shot diagnosis of underrepresented abnormalities.
{IEEEkeywords}
Computed tomography (CT), cross-modal interaction, vision-language alignment, zero-shot learning
1 Introduction
Figure 1: UMAP visualization of features. Cosine similarity is used to evaluate the alignment between image and text features. The text features are generated using generic descriptive texts to ensure that all images can be matched to all texts. (a) Features generated using pretrained weights without vision-language alignment, where image (in blue) and text features (in red) remain unaligned. The corresponding pre-training protocols are detailed in “Implementation Details”. (b) Features after vision-language alignment using 3D CLIP, showing improved alignment but with noticeable modality gaps. (c) Features after vision-language alignment using BrgSA framework, where the CMKB features (in green) serve as a bridge to reduce the modality gap and further enhance feature alignment. CMKB denotes the Cross-Modal Knowledge Bank, whereas BrgSA abbreviates Bridged Semantic Alignment.
@@ 1 Introduction
Computed Tomography (CT) is a widely used non-invasive diagnostic tool in clinical practice, playing a crucial role in detecting various abnormalities [1]. With the rapid development of artificial intelligence, significant progress has been made in leveraging CT images for automated abnormality diagnosis [2]. Supervised learning methods have achieved impressive performance in tasks such as disease classification and detection [3, 4]. However, supervised learning methods heavily rely on large-scale, manually annotated datasets, which is time-consuming and labor-intensive [5]. Moreover, the reliance on extensive annotations limits the diversity of available training data, particularly for rare diseases, where labeled samples are scarce [6]. This constraint hampers the scalability of supervised learning methods and diminishes their ability to generalize across a wide spectrum of diseases [7].
@@
Recently, zero-shot learning (ZSL) based on vision-language alignment has emerged as a promising alternative to traditional supervised learning methods [8, 9]. By eliminating the need for extensive manual annotations, ZSL offers a more efficient path to diagnose a diverse range of abnormalities. For instance, CLIP [10] leverages contrastive learning on large-scale image-text datasets to enable effective vision-language alignment and perform zero-shot classification without extensive manual annotations [11]. In the medical domain, several studies have demonstrated the potential of CLIP for zero-shot abnormality diagnosis on 3D CT volumes [12, 13, 14, 15]. However, directly applying CLIP to 3D medical images remains inadequate for vision–language alignment due to a significant modality gap between image and text embeddings. As shown in Fig. 1(a,b), image and text features form two distinct clusters even after alignment, with the feature alignment score (FAS) improving only from 0.08 to 0.37. This persistent separation indicates that standard CLIP fails to fully capture the semantic correspondence between visual and textual representations in the medical domain. Similar modality discrepancies have also been observed in previous studies [16, 17]. This highlights the necessity of addressing the modality gap in CLIP embeddings to achieve better alignment in 3D medical images.
@@
In this paper, we propose a simple yet effective framework named Bridged Semantic Alignment (BrgSA), which aims to bridge the modality gap and facilitate effective vision-language alignment. Our approach consists of two modules: semantic summarization and cross-modal knowledge interaction (CMKI). First, we utilize a large language model (LLM), which has powerful semantic understanding capabilities, to summarize clinical reports and extract critical information. This significantly reduces the learning difficulty on the textual side and produces high-level semantic features. Next, we propose a CMKI module to supplement the conventional CLIP alignment, where a cross-modal knowledge bank (CMKB) acts as a semantic bridge between image and text features. By projecting image and text features into the shared latent space of the CMKB, it reduces the discrepancy between modalities and simultaneously preserves their unique characteristics. This enables an implicit alignment approach that does not rely on paired image-text data. Finally, we employ contrastive learning to compare positive and negative pairs, achieving explicit vision-language alignment. Fig. 1(c) visualizes CMKB serving as a bridge between image and text features, effectively narrowing the modality gap. Also, the FAS is further improved to 0.50, indicating a better alignment.
@@
To explore the potential of our vision-language alignment approach for zero-shot diagnosis, we expand the original 18 abnormality labels on CT-RATE [12] with 15 additional labels. The new benchmark, referred to as “CT-RATE-LT”, focuses on underrepresented abnormalities with limited occurrences in the dataset, providing a challenging yet valuable testing ground for evaluating diagnostic methods. Experimental results on CT-RATE-LT demonstrate that our method achieves excellent performance in diagnosing underrepresented abnormalities, showing significant improvements compared to the existing state-of-the-art (SOTA) method (AUC: 76.9 vs. 85.6). Furthermore, we evaluate our method on the benchmark CT-RATE [12] and RAD-ChestCT [18] datasets, achieving SOTA performances in zero-shot abnormality diagnosis, with an AUC of 79.2 and 70.0 on internal and external validation, respectively. In addition to its diagnostic capabilities, our method significantly outperforms the SOTA in the report-to-volume retrieval task (Recall@10: 5.0 vs. 10.1), demonstrating effective semantic alignment between image and text features.
@@ 2.1 Medical Vision-language Pretraining
Existing medical vision-language pretraining (VLP) research is predominantly focused on 2D imaging. Public datasets like MIMIC-CXR [19] have provided a solid foundation for the advancement of 2D VLP, while PadChest [20], which encompasses 193 abnormality categories, has established an effective benchmark for rare disease diagnosis. GLoRIA [21] enhances the ability to capture associations between images and text by integrating global and local feature alignment. CheXZero [22] leverages a CLIP model pre-trained on natural data, achieving stable improvements in medical VLP performance. MedCLIP [23] employs unpaired medical images and text with a semantic matching loss to mitigate false negatives. Xplainer [24] introduces an explainable zero-shot diagnosis framework based on observation-driven contrastive learning. MedKLIP [25] and KAD [26] incorporate domain-specific medical knowledge to improve VLP performance in CXR diagnosis tasks. CARZero [27] advances cross-modal alignment by leveraging cross-attention mechanisms to address the complex relationships between visual and textual modalities. These advancements have established 2D VLP as a robust framework for various abnormalities diagnosis.
@@ 2.2 Multi-Modality Interaction
Cross-modal interaction has been extensively explored to bridge the gaps between modalities, such as vision and language modalities. Transformer-based approaches like ViLBERT [29] employs dual-stream transformers with co-attentional layers for modality-specific processing, while ViLT [30] uses a unified transformer to jointly encode visual patches and textual embeddings, enabling efficient cross-modal interaction. MGCA [31] introduces cross-attention mechanisms to achieve cross-modal interaction, aligning medical image and text representations at multiple levels. MPMA [32] enhances cross-modal interaction by integrating image-text reconstruction with a global and local alignment mechanism, enabling richer semantic representation learning. However, these approaches primarily emphasize explicit alignment mechanisms, often overlooking the potential of leveraging shared semantic spaces to guide cross-modal learning.
Dictionary learning has emerged as a technique for enhancing cross-modal alignment by constructing shared representational spaces. Deng et al. [33] aligns features via a shared label space to improve cross-modal retrieval. HCDDL [34] builds hierarchical semantic embeddings by dictionary learning for fine-grained alignment. UNIMO-2 [35] leverages grounded semantic spaces to bridge visual and textual modalities, effectively addressing misalignment. Methods such as DetCLIP [36] and SOHO [37] further exploit trainable dictionaries to enrich visual-semantic representations, achieving SOTA results in zero-shot and vision-language tasks. CP-CLIP [38] employs the Core–Periphery principle to facilitate cross-modal interaction by structuring shared semantics and modality-specific features within a unified latent space. Despite these advances, prior dictionary-based methods mainly rely on reconstruction losses, lacking discriminative constraints and risking cross-modal leakage. To overcome these issues, we design a Cross-Modal Knowledge Interaction (CMKI) module that jointly optimizes reconstruction and contrastive objectives within separate modality-specific banks. This implicit–explicit synergy preserves modality integrity and ensures robust alignment for 3D medical imaging.
Figure 2: Illustration of the proposed BrgSA network, which integrates semantic summarization and cross-modal knowledge interaction (CMKI). First, we leverage a large language model (LLM) to summarize the report, generating outputs in a fixed template. These summarized reports, along with the original reports, serve as the textual inputs. Then, image and text features are extracted by respective encoders and fed into CMKI module to obtain interaction features. Finally, the interaction features are constrained using an MSE loss, while alignment optimization is achieved via an InfoNCE loss.
@@ 3 Method
In this section, we present the proposed BrgSA framework for zero-shot classification, consisting of semantic summarization and a cross-modal knowledge interaction (CMKI) module (Fig. 2). Further details are provided below.
@@ 3.1 Semantic Summarization
To simplify learning on the textual side, we introduce semantic summarization. In the context of medical reports, textual data often contains lengthy and complex descriptions, including both critical abnormality-related information and a significant amount of irrelevant details. Such data characteristics increase the difficulty of extracting and aligning core semantic information with image features. We leverage the powerful semantic understanding and summarization capabilities of LLM to produce semantic summaries of medical reports, which are much easier to understand. As illustrated in Fig. 3, a prompt is designed to guide the LLM in extracting key abnormality-related information and generating outputs in a fixed template: “There is [abnormality].” This template-based summarization significantly reduces the alignment difficulty by providing more focused and consistent descriptions.
However, solely relying on summarized reports may lead to information loss, especially when subtle or nuanced abnormalities are omitted. To balance simplicity and completeness, we propose a dual-input strategy that incorporates both the original report and its corresponding summary. This combined representation retains the rich contextual information from the full report while benefiting from the semantic clarity of the summary. As a result, it enhances the robustness and accuracy of image-text alignment in BrgSA training.
Figure 3: Prompt for LLM used in semantic summarization for reports.
@@ 3.2 Feature Extraction
Assume that the training dataset contains $N$ samples denoted as $S_{\text{train}}=\left\{\left(I_{1},R_{1}\right),\ldots,\left(I_{N},R_{N}\right)\right\}$, where $I_{i}\in\mathbb{R}^{H\times W\times D}$ represents a 3D CT volume and $R_{i}$ represents its corresponding medical report. Specifically, $R_{i}$ refers to the concatenation of the original medical report and its semantic summarization, which is the dual-input strategy. The combined text is then used for alignment with the image features. Here, $H,W,D$ denote the height, width, and depth of the 3D CT volumes, respectively. As illustrated in Fig. 2, we introduce the components of our architecture for feature extraction, including an image encoder $\mathbf{\Phi}_{\text{I}}$ for CT images and a text encoder $\mathbf{\Phi}_{\text{R}}$ for medical reports.
3.2.1 Image Encoder
The image encoder is designed to extract global features from 3D CT images, as shown in Eq. 1.
@@ 3.2.1 Image Encoder
where $V_{i}\in\mathbb{R}^{d}$ represents the global features of the image, and $d$ denotes the feature dimension. In our experiments, we adopt a ViT-base [39] model as the image encoder. The output of the [CLS] token is used to represent the global feature for 3D image, following the methodology in [39].
3.2.2 Text Encoder
The text encoder is used to extract global features from the medical report, as shown in Eq. 2.
@@ 3.2.2 Text Encoder
where $T_{i}\in\mathbb{R}^{d}$ represents the global features of the medical report. We use CXRBERT [40] as the text encoder, which is specifically pretrained on large-scale radiology reports and thus better suited for chest imaging tasks. The output of the [CLS] token is used as the global feature for the text [41]. In our experiment, the output feature dimensions of both the image and text encoders are unified to $d$.
3.3 Cross-Modal Knowledge Interaction
@@
In the CMKI module, we introduce a cross-modal knowledge bank (CMKB) as a bridge for image-text interaction. The CMKB is initialized as a set of learnable embeddings:
@@
where each $\mathbf{b}_{k}$ denotes a basis vector and $K$ represents the number of basis vectors.
@@
Using the attention weights $\mathbf{Z}_{i}^{V}$ and $\mathbf{Z}_{i}^{T}$, the original image and text features are represented in the CMKB space by taking a weighted combination of the basis vectors, yielding their approximated representations:
@@
where $\alpha$, $\beta$, and $\gamma$ are hyperparameters controlling the contributions of the reconstruction loss ($\ell_{\text{MSE}}$), the InfoNCE loss ($\ell_{\text{INFO}}$), and the InfoNCE loss on reconstructed features ($\ell_{\text{INFO-R}}$), respectively. To further clarify the computational flow, we provide a concise pseudocode implementation of the CMKI module in Algorithm 1.
Algorithm 1 Cross-Modal Knowledge Interaction
1:Image features $V_{i}$, text features $T_{i}$, CMKB $B$, temperature $\tau$
[…]
Figure 4: Histogram of abnormality frequencies for CT-RATE-LT.
Figure 5: Mapping of 27 abnormalities from RAD-ChestCT to 18 abnormalities in CT-RATE. The abnormalities from CT-RATE are denoted in blue font, whereas the abnormalities from RAD-ChestCT are denoted in black font.